OpenStudy (anonymous):

log equation in exponential form: log10 1/1000 =-3

OpenStudy (dumbcow):

Here is the general forms for equivalent log and exponential forms where b is base logb (x) = y <==> x = b^y so your base is 10, x = 1/1000, y = -3 ->1/1000 = 10^-3 make sense??

OpenStudy (anonymous):

yesss thanks :)

OpenStudy (dumbcow):

your welcome